OpenStudy (thai23):

2|3x-1|+2=10 answers A. X= -11/6 or X=3/2 B. X= 3/2 or X=5/3 C. X=3/2 or X= -1 D. X= -1 or X= 5/3

OpenStudy (ben00):

first do 2 times 3x and 2 times -1.

OpenStudy (thai23):

okay

OpenStudy (ben00):

6x - 3 i guess?|

OpenStudy (thai23):

i don't think so? lol

OpenStudy (ben00):

Its really hard. I think I am also learning this

OpenStudy (ben00):

2 x 3x would = 3x and 2 x -1 = -2

OpenStudy (ben00):

So maybe after that 6x -2 + 2 = 10 6x = 10

OpenStudy (ben00):

Then do - i think 6x - 6 = 10 - 6 =

OpenStudy (ben00):

x = 4?

Awolflover1 (awolflover1):

Add -2 to both sides. \[2(|3x−1|)+2+−2=10+−2\]

Awolflover1 (awolflover1):

\[10+-2=? \]